40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 25 16:56:44 2022, Paco, Configuration, BHD, IFO recovery - IMC alignment 
    Reply  Wed May 25 17:34:47 2022, yuta, Configuration, BHD, IFO recovery - IFO alignment Screenshot_2022-05-25_17-47-57.png
       Reply  Thu May 26 19:55:43 2022, yuta, Configuration, BHD, Oplevs centered, BHD DCPDs are now online elog_1Y2.JPGelog_BHD.JPGelog_box.JPGScreenshot_2022-05-26_17-37-27_IFOaligned_OplevCentered.pngScreenshot_2022-05-26_20-35-02.png
          Reply  Fri May 27 17:45:53 2022, yuta, Configuration, BHD, BHD camera installed, GRY aligned 
             Reply  Thu Jun 2 20:05:37 2022, yuta, Configuration, PSL, IMC input power recovered to 1W, some alignment works Before.JPGAfter.JPG
                Reply  Fri Jun 3 12:13:58 2022, Paco, Configuration, CDS, Fix RFM channels  SoGreen.png
                   Reply  Fri Jun 3 15:22:51 2022, yuta, Update, LSC, Both arms locked with POY/POX, IR beam centered on TMs with ASS IRBeamsOnTMs.JPGScreenshot_2022-06-03_15-03-51.png
                      Reply  Fri Jun 3 17:42:50 2022, yuta, Update, LSC, MICH locks with AS55_Q Screenshot_2022-06-03_17-41-55.png
                         Reply  Sun Jun 5 19:46:40 2022, Paco, Update, LSC, Fixed IMC Trans sum issue 
                         Reply  Mon Jun 6 13:35:11 2022, Paco, Update, LSC, First calibrated spectra of MICH at AS55 Q Screenshot_2022-06-06_17-30-16_MICHOLTF.pngCalibrated_MICH_ERR.pdf
                            Reply  Fri Jun 17 16:22:21 2022, yuta, Update, LSC, Actuator calibration of BS. ITMX, ITMY, updated MICH displacement spectra from c1cal 6x
                               Reply  Thu Jul 7 18:18:19 2022, yuta, Update, LSC, Actuator calibration of ETMX and ETMX Screenshot_2022-07-05_14-52-01_OLTF.pngScreenshot_2022-07-05_14-54-03_TF.pngScreenshot_2022-07-05_14-56-41_Ratio.png
                                  Reply  Thu Jul 7 18:22:12 2022, yuta, Update, LSC, Actuator calibration of MC2 using Yarm TF.pngMC2.png
                                     Reply  Fri Jul 8 16:18:35 2022, rana, Update, LSC, Actuator calibration of MC2 using Yarm 
                                     Reply  Mon Jul 18 17:07:12 2022, yuta, Update, LSC, x4.12 added to ETMX coil outputs to balance with ETMY 
          Reply  Thu Jun 16 18:23:15 2022, Paco, Configuration, BHD, Recovering LO beam in BHD DCPDs Screenshot_2022-06-16_18-29-14_BHDLObeamISBACK.png
          Reply  Tue Jun 21 14:17:50 2022, yuta, Configuration, BHD, BHD DCPDs re-routed to c1sus2 C1X07ADC1.JPGBHDDCPDs.JPG
             Reply  Tue Jul 19 16:00:34 2022, yuta, Configuration, BHD, Fast channels for BHD DCPDs now available in c1lsc but not in c1hpc Screenshot_2022-07-19_14-26-39_c1hpc.pngScreenshot_2022-07-19_14-24-49_c1lsc.pngScreenshot_2022-07-19_15-51-25_GreenGreen.png
                Reply  Thu Jul 21 21:50:47 2022, Tega, Configuration, BHD, c1sus2 IPC update 
                   Reply  Fri Jul 22 15:05:26 2022, Tega, Configuration, BHD, c1sus2 shared memory and ADC fix Screenshot_2022-07-22_15-04-33_DCPD.pngScreenshot_2022-07-22_15-12-19_models.pngScreenshot_2022-07-22_15-15-11_ERR.pngScreenshot_2022-07-22_15-32-19_GDS.png
                      Reply  Fri Jul 22 17:46:10 2022, yuta, Configuration, BHD, c1sus2 watchdog update and DCPD ERR channels Screenshot_2022-07-22_17-48-25.png
                         Reply  Mon Jul 25 17:58:10 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update Screen_Shot_2022-07-25_at_5.43.28_PM.pngScreen_Shot_2022-07-25_at_5.21.10_PM.pngScreen_Shot_2022-07-25_at_5.30.58_PM.pngScreen_Shot_2022-07-25_at_5.35.19_PM.png
                            Reply  Mon Aug 1 18:42:39 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update c1sus2_dolphin.pngfb1_dxamin_status.pngdolphin_num_mem_init2.png
                               Reply  Tue Aug 2 17:25:18 2022, Tega, Configuration, BHD, c1sus2 dolphin IPC issue solved c1lsc_IPC_status.pngFE_lsmod_dependencies_c1sus2_b4_after_iop_unpdate.png
                      Reply  Mon Jul 25 18:09:41 2022, Tega, Configuration, BHD, BHD Homodyne Phase control MEDM screen Screen_Shot_2022-07-25_at_6.12.08_PM.pngScreen_Shot_2022-07-25_at_6.18.09_PM.png
Message ID: 16886     Entry time: Thu Jun 2 20:05:37 2022     In reply to: 16880     Reply to this: 16887
Author: yuta 
Type: Configuration 
Category: PSL 
Subject: IMC input power recovered to 1W, some alignment works 

[Paco, Yuta]

We have increased the output power from the PSL table to 951 mW (it was 96.7 mW).
IMC was recovered including WFS, and both arms are flashing nicely in IR.
We tweaked the alignment of GRX and GRY injection to align them with IR, but it was hard.
Right now IR beams are not centered on TMs. We should center them first.

What we did:
Power increase and IMC recovery
 - Replaced a beam splitter which splits the beam into IMC REFL RF PD path and WFS path from R=98% to R=10% one. Reflection goes to RF PD.
 - Put a R=98% beam splitter back into WFS path.
 - We also tried to put a window in front of IMC REFL camera to recover the arrangement in 40m wiki, but the beam reflected from the window was too weak for us to align. So, we decided not to place a window in front of the camera.
 - Attached photos are the IMC REFL path before and after the work.
 - Measured the PSL output power as Koji did in elog 40m/16672. It was measured to be 96.7+/- 0.5 mW.
 - Rotated the HWP using the Universal Motion Controller (it was not possible for us to do it from the MEDM screen). The position was changed from 73.99 deg to 36.99 deg. Output power was measured to be 951 +/- 1 mW
 - IMC locked without any other changes.
 - Changed C1:IOO-WFS_TRIGGER_THRESH_ON to 5000 (was 500). IMC WFS also worked.
 - After running MC WFS relief script, WFS DC offsets and RF offsets are adjusted following the steps in elog 40m/16835. Below are the results.

C1:IOO-WFS1_SEG1_DC.AOFF => -0.0008882080010759334
C1:IOO-WFS1_SEG2_DC.AOFF => -0.0006527877490346629
C1:IOO-WFS1_SEG3_DC.AOFF => -0.0005847311617496113
C1:IOO-WFS1_SEG4_DC.AOFF => -0.0010395992663688955
C1:IOO-WFS2_SEG1_DC.AOFF => -0.0025944841559976334
C1:IOO-WFS2_SEG2_DC.AOFF => -0.003191715502180159
C1:IOO-WFS2_SEG3_DC.AOFF => -0.0036688060499727726
C1:IOO-WFS2_SEG4_DC.AOFF => -0.004011172490815322


IOO-WFS1_I1         :  +1977.7 ->    +2250 (Significant change)
IOO-WFS1_I2         :  +3785.8 ->  +3973.2
IOO-WFS1_I3         :  +2014.2 ->  +2277.7 (Significant change)
IOO-WFS1_I4         :  -208.83 ->  +430.96 (Significant change)
IOO-WFS1_Q1         :  +2379.5 ->  +1517.4 (Significant change)
IOO-WFS1_Q2         :  +2260.4 ->  +2172.6
IOO-WFS1_Q3         :  +588.86 ->  +978.98 (Significant change)
IOO-WFS1_Q4         :  +1654.8 ->  +195.38 (Significant change)
IOO-WFS2_I1         :  -1619.9 ->  -534.25 (Significant change)
IOO-WFS2_I2         :  +1610.4 ->  +1619.8
IOO-WFS2_I3         :  +1919.6 ->  +2179.8 (Significant change)
IOO-WFS2_I4         :    +1557 ->  +1426.6
IOO-WFS2_Q1         :   -62.58 ->  +345.56 (Significant change)
IOO-WFS2_Q2         :  +777.01 ->  +805.41
IOO-WFS2_Q3         :  -6183.6 ->  -5365.8 (Significant change)
IOO-WFS2_Q4         :  +4457.2 ->  +4397.


IFO Alignment
 - Aligned both arms using IR. Both arm flashes at the following, which is consistent with the power increase.
 C1:SUS-ETMX_TRX_OUT_DQ ~1.1
 C1:SUS-ETMY_TRY_OUT_DQ ~0.6
 - With this, we tried to tweak GRX and GRY injection. The following is after the work. We could increase GTRX to 0.204 when the Xarm is aligned to green. This suggests that GRX injection is not aligned nicely yet. But the beams are also not centered on TMs. We should center them first.
 C1:ALS-TRX_OUT_DQ ~0.13
 C1:ALS-TRY_OUT_DQ ~0.07
 - GTRX and GTRY cameras are adjusted to have nicer images. In GRX path, the second and last lens before the PD and CCD was pulled ~ 1 cm behind its original position and both beams realigned. Then, on GRY path, the beam was re-centered on the first and only lens, the whole assembly pushed forward by ~ 2 cm and the beams re-centered.

Next:
 - Center the IR beam on TMs (first by our eyeballs; better to use A2L after arm locking is recovered and coils are balanced)
 - Tweak GRX and GRY injection (restore GRY PZTs?)
 - Install ETMXT camera (if it is easy)
 - Lock Xarm and Yarm (C1:LSC-TRX/Y_OUT needs to be fixed for triggering. Can we use other PDs for triggering?)
 - MICH locking (REFL and AS PDs might need to be re-aligned; they are not receiving much light)
 - RTS model for BHD needs to be updated

Attachment 1: Before.JPG  4.556 MB  Uploaded Thu Jun 2 21:44:29 2022  | Hide | Hide all
Before.JPG
Attachment 2: After.JPG  4.763 MB  Uploaded Thu Jun 2 21:44:43 2022  | Hide | Hide all
After.JPG
ELOG V3.1.3-